OpenStudy (anonymous):

Can someone help me and explain how to solve this system of equation? Will medal and fan! x+y=5 .10x+.25y=20

OpenStudy (welshfella):

you can solve this by substitution from the first equation:- x = 5 - y

OpenStudy (welshfella):

so replace the x in the second equation by 5 - y and solve this equation for y

OpenStudy (anonymous):

Elimination would be easier.

OpenStudy (anonymous):

How would elimination be easir?

OpenStudy (anonymous):

*easier?

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

Nvm. :\

OpenStudy (anonymous):

Okay, @welshfella thescond equation with the combination method would be .10(5-y)+.25y=20

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Okay so .5-.5y+25y=20 .5-30y=20

OpenStudy (anonymous):

I mean. 25.5y

OpenStudy (welshfella):

isn't it 0.25y? 5 - .10y + .25y = 20 .15y = 15

OpenStudy (anonymous):

Okay, so y=100?

OpenStudy (welshfella):

No - i made a mistake in my last post its 0.5 not 5 so we have 0.15y = 20 - 0.5 = 19.5

OpenStudy (anonymous):

130?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

no plug in 130 to x + y = 5 to get value of x

OpenStudy (welshfella):

*now

OpenStudy (anonymous):

x+130=5?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Okay, I think I have it from here, thanks!

OpenStudy (welshfella):

yw